OpenStudy (anonymous):

WILL AWARD MEDAL The lengths of two sides of a triangle are shown below: Side 1: 8x2 - 5x - 2 Side 2: 7x - x2 + 3 The perimeter of the triangle is 4x3 - 3x2 + 2x - 6. Part A: What is the total length of the two sides, 1 and 2, of the triangle? Part B: What is the length of the third side of the triangle? Part C: Do the answers for Part A and Part B show that the polynomials are closed under addition and subtraction? Justify your answer

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@kirbykirby

OpenStudy (anonymous):

(11x-2)(5x+3)?

OpenStudy (anonymous):

im guessing i have no idea how to solve this -.-.....

OpenStudy (kirbykirby):

Part A: just add the 2 sides together (Side 1) + (side 2) = ?

OpenStudy (anonymous):

how exactly? i need a walkthrough...because the answers are probably going to be wrong.

OpenStudy (anonymous):

*the answers im getting are probably

OpenStudy (kirbykirby):

\( (8x^2 - 5x - 2)+ (7x - x^2 + 3) \) Now just add up like-terms (add up the \(x^2\)-terms together, and add the \(x\)-terms together

OpenStudy (anonymous):

7x^2+2x+1

OpenStudy (kirbykirby):

yes :)

OpenStudy (kirbykirby):

Now for part b), realize that the perimeter is the sum of all 3 sides of the triangle.. so (side 1) + (side 2) + (side 3) = \(\large 4x^3 - 3x^2 + 2x - 6\) \((7x^2 + 2x + 1)+(\text {side 3})=4x^3-3x^2+2x-6\)

OpenStudy (kirbykirby):

Isolate for (side 3)... so you subtract the perimeter from the sum of the 2 other sides

OpenStudy (anonymous):

so its 4x^3-10x^2-7.THANKS!